A car is moving at 17.7 m/s when it begins to slow down with a deceleration of 2.5 m/s2. How long does it take for the car to travel 80 meters as it slows down?

 


1
Expert's answer
2021-09-22T00:54:15-0400
"v(t)=\int adt=v_0+at"

"s(t)=\int vdt=\int (v_0+at)dt=s_0+v_0t+\dfrac{at^2}{2}"

Given "a=-2.5\ m/s^2, v_0=17.7\ m/s,s _0=0\ m"


"v(t_1)=0=>17.7-2.5t_1=0"

"t_1=7.08 \ s"

"s(t_1)=s(7.08)=0+17.7(7.08)-\dfrac{2.5(7.08)^2}{2}"

"=62.658 (m)"

Car travels 62.658 meters and stop.

Therefore car cannot travel 80 meters.